OpenStudy (haileeey.brooke):

How many days would it take a construction loan of $548,048 to earn$50,000 exact interest at 9% interest rate

OpenStudy (shadowlegendx):

Do you recognize the term, principal amount?

OpenStudy (haileeey.brooke):

yes

OpenStudy (shadowlegendx):

Do you know what it means?

OpenStudy (haileeey.brooke):

the amount borrowed separate from interest

OpenStudy (shadowlegendx):

What would that be here?

OpenStudy (haileeey.brooke):

548,048 because if the interest is 50,000 it couldnt be that and its not the 9% rate correct?

OpenStudy (shadowlegendx):

A loan is someone borrowing money

OpenStudy (shadowlegendx):

And $548,048 was taken as a construction loan, so :)

OpenStudy (haileeey.brooke):

so you take the interest rate of 9% but what im not getting is how you find how many days it will take :/

OpenStudy (shadowlegendx):

Did your lesson or teacher every provide a formula for you?

OpenStudy (shadowlegendx):

Also please list any multiple choice answers, if any

OpenStudy (haileeey.brooke):

370 days, 360 days, 365 days, and 364 days

OpenStudy (shadowlegendx):

Well, what we're trying to solve for, is t, or the amount of time it would take to obtain $50,000 in interest, from a construction loan of $548,048 at a 9% interest rate

OpenStudy (shadowlegendx):

So here is the formula I will be using Interest Amount = (Principle Amount)(Interest Rate)t/365 Lets make that prettier Let I = Interest Amount Let P = Principal Amount Let R = Interest Rate Let t = time \[I = (PRt)/365\]

OpenStudy (shadowlegendx):

Can you solve it now?

OpenStudy (haileeey.brooke):

yes i think so, so the I=50,000 and the P=548,048 and the R= 9% and then we're trying to find time?

OpenStudy (shadowlegendx):

time basically = days

OpenStudy (shadowlegendx):

Only in this equation though, as I have that side of the equation being divided by 365 which converts it from a year, to days

OpenStudy (haileeey.brooke):

oh, i see now so you plug in 365 for time?

OpenStudy (shadowlegendx):

No

OpenStudy (haileeey.brooke):

im so confused do you leave it as blank and solve for it?

OpenStudy (shadowlegendx):

Yes leave t blank. We're solving for t

OpenStudy (shadowlegendx):

Like 3x = 15 "Solve for x"

OpenStudy (haileeey.brooke):

and then t is the final answer?

OpenStudy (shadowlegendx):

Just looks a bit different :)

OpenStudy (shadowlegendx):

yes

OpenStudy (shadowlegendx):

Let me explain it all in one big post for you

OpenStudy (haileeey.brooke):

okay thank you!!

OpenStudy (shadowlegendx):

Ok forget anything I've said before. Problem: "How many days would it take a construction loan of $548,048 to earn$50,000 exact interest at 9% interest rate" So what are we trying to find? The amount of days it will take an interest rate of 9% on $548,048 to make exactly $50,000. So now we setup the equation, to solve for the amount of days it would take. Let t = the amount of days it would take We start with 50,000 as we know that at some point, t multiplied by the $548,048, multiplied by the interest rate, then multiplied t, all over 365, will get -> 50000 So 50,000 = (548,048)(0.09)t/365 Interest Amount = (Principle Amount)(Interest Rate)t/365

OpenStudy (shadowlegendx):

I spent a lot of time trying to figure out an easier formula for you, so that you could remember it for future use but it got con-jumbled o-o Just follow the above

OpenStudy (shadowlegendx):

Just setup this problem So 50,000 = (548,048)(0.09)t/36 and solve for t using algebra

OpenStudy (shadowlegendx):

So 50,000 = (548,048)(0.09)t/365